# Zhang, Liu, and Zhao

### v4

@@ 2.3 Consistency comparison with existing DR-based methods
[…]
In comparison, our PADR-based ERM model is nonasymptotically consistent in more general cases, including convex SP problems and nonconvex ones. Moreover, by Corollary 2.4, if there exists a convex or DC optimal DR, our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large. This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $[-M,M]^{d}$ by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
@@ 3 Efficient Sampling-based Algorithm for Learning PADR
where the objective is the finite sum of composite-PA functions. Instead of attempting for global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions. Among various kinds of stationary concepts, the d-stationary solution is the sharpest type of first-order stationary solutions for general nonconvex optimization. For the PADR learning problem (9) with a general nonconvex and nondifferentiable composite structure, our algorithmic development is mainly motivated by the following question:
@@
Building on this motivation, we develop an MM-type algorithm specializing to the structured PADR-based ERM problem with the theoretical convergence analysis. It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which speeds up to 4 times faster than the basic MM method.
To make our discussion self-contained, we first present the standard MM algorithms in Section 3.1, which serves as a reference for the motivation behind our proposed sampling-based algorithm in Section 3.2. We make some remarks on notations to facilitate algorithmic development and convergence analysis. To emphasize the role of PADR as a function of $\theta$, with $\xi=(\xi_{x},\xi_{y})\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,\xi_{x})$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,\xi_{x})$ where $\{g_{i}(\theta,\xi_{x})\}$ and $\{h_{i}(\theta,\xi_{x})\}$ are linear functions with respect to $\theta$ and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
@@ 3.1 Preliminary analysis: standard MM algorithms
A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021, Definition 7.1.1). Under (P1) and (P2), it yields the sequence of decreasing objective values as follows,
@@
Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by state-of-the-art convex-programming solvers. By the basic analysis (see Chapter 7 in Cui and Pang (2021)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$. Such a fixed point property could be related to various types of stationarity of the original problem depending on the surrogate functions as illustrated in Cui et al. (2018) and Liu et al. (2022).
Regarding the PADR-based ERM problem (10), due to the nonconvex and nondifferentiable piecewise affine functions, a major issue of the standard MM algorithms including Lipp and Boyd (2016) and Le Thi and Pham Dinh (2018) is that the fixed point corresponding to the surrogation map is a (compound) critical solution (see definitions in Pang et al. (2016), Liu et al. (2022)) which is much weaker than directional stationarity. Specifically, based on the composite structure in (10), approaches to construct surrogate functions of $F(\theta,\xi)$ in Cui et al. (2018) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
@@
They are the upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$ respectively. For $d>1$, the maximum is taken componentwisely. We illustrate the construction of the surrogate function when outer function $\varphi$ is linear as an example. With $\varphi(z,\xi)=\xi^{\top}z$, we construct the surrogate function of $F(\cdot,\xi)$ as $\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{\xi,0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{\xi,0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$, where $I\coloneqq\big{(}I_{1},I_{2}\big{)}\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$ for any $\xi\in\Xi$. It obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm. Clearly $f^{\prime}(\cdot,\xi)(\theta;d)=g^{\prime}(\cdot,\xi)(\theta;d)-\max\{\,\nabla h_{i}(\bar{\theta},\xi)^{\top}d:\,i\in\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi)\}$, whereas the surrogate function $\widehat{F}(\theta,\xi;\bar{\theta},I)$ employs an arbitrary active index $I\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$. Hence, the fixed point of the surrogation map is a critical solution. Solutions of criticality are not desirable because in practice solutions may be at the “cliff” of the cost landscape, which is illustrated in Example 3.1 and Figure 1b.
@@
Consider problem (10) with $d=2$, $q=1$, $\Xi=\{\xi_{0}\}$, $\varphi(z,\xi_{0})=z_{1}^{2}+z_{2},f_{1}(\theta,\xi_{0})=\theta$, and $f_{2}(\theta,\xi_{0})=-\max\{-\theta,0\}$. Figure 1a illustrates the two candidates of surrogate functions depending on which piece of the inner PA function $f_{2}(\theta,\xi_{0})$ is active at the reference point of interest. Figure 1b and Figure 1c illustrate the iterative procedures of the standard MM algorithm and the enhanced MM algorithm respectively, with the initial point $\theta_{0}=1.1$. The standard MM algorithm converges to a critical point $\theta_{\infty}=0$. For the enhanced MM algorithm with $\varepsilon=0.3$, at the second iterate point $\theta_{2}$, we have $\left\lvert\mathcal{I}^{\varepsilon}(\theta_{2},\xi_{0})\right\rvert=2$ at $\theta_{2}$, which thus leads to the next iterate point passing over 0 and continues to move towards a d-stationary point $\theta^{d}$ that is also globally optimal to the problem.
@@
Despite possessing stronger convergence properties, the enhanced MM algorithm applied for large-scale ERM problems could be exceptionally inefficient as it requires iteratively solving a potentially very large number of finite-sum convex subproblems over all the data. Sampling-based MM-type algorithms have been recently studied for (partially) smooth stochastic DC programs in Nitanda and Suzuki (2017), Xu et al. (2019), and Le Thi et al. (2022) and nondifferentiable stochastic composite DC problems in Liu et al. (2022) to compute compound critical solutions. Based on a natural idea of combining a sequential sampling scheme with the enhanced technique, we propose an enhanced sampling-based MM algorithm with rigorous convergence analysis in the next subsection to solve the stochastic composite DC programs (10) with both computational efficiency and theoretical convergence guarantee.
@@ 3.2 Enhanced sampling-based MM algorithm for learning PADR
The directional derivative consistency condition is introduced in Cui and Pang (2021) to relate the directional derivatives of surrogation and original functions. We consider the following three cases of the outer function $\varphi$ covering a broad class of cost functions.
@@
For the above cases, we can construct the surrogate function of $F(\cdot,\xi)$ satisfying Assumptions (B1) to (B4), by utilizing surrogate functions (12) and (13) of the inner PA function $f(\cdot,\xi)$ with the selected index $I=(I_{1},I_{2})$. The specific construction and verification are provided in Online Appendix 7. In particular, Case 2 is a multivariate extension of the one considered in Cui et al. (2018). Case 3 is suitable for nonconvex smooth cost functions that are applicable for many practical problems.
@@
For a general nonconvex problem on a convex set with the directional differentiable objective, d-stationarity is identified as the sharpest kind among the first-order stationarity concepts. Specialized to the composite DC program (10), according to Cui et al. (2018, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2. Lu et al. (2019) proposes the strengthened d-stationarity for structured DC programs which takes into account the $\epsilon$-active index set. We extend to the concept of composite $(\varepsilon,\rho)$-strong d-stationarity specialized for the PADR-based ERM problem (9). The composite strongly d-stationary point can be interpreted as the fixed point corresponding to the proximal mapping of the expected minimum surrogate function $\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};\bar{\theta})]$ based on the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfying Assumptions (B1)-(B4).
@@
Due to the requirement of a unique proximal mapping point in the defining condition, our concept is slightly stronger than $(\alpha,\eta)$-d-stationarity defined in Lu et al. (2019) when the problem is reduced to composite DC program. Moreover, it is weaker than composite $\varepsilon$-strong d-stationarity introduced in Qi et al. (2021b) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis. Furthermore, in Online Appendix 8.1, we show that $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ under a regularity assumption on $F$. It implies that the definition of $\mathcal{S}^{d}_{\varepsilon,\rho}$ serves as a natural connection between d-stationarity and global optimality.
@@
We propose the ESMM algorithm with a natural idea of combining the randomized $\varepsilon$-active surrogation with sequential sampling presented in Algorithm 1. At iteration $\nu$, a subset of samples $\Xi_{\nu}=\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ of size $N_{\nu}$ is generated in step 2. Then in step 3 a sampling-based surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ is constructed which can be regarded as the upper surrogation function to the sample average approximation function $F_{\nu}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot,\widetilde{\xi}_{\nu}^{s})$ with the combination of indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}=\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ randomly selected from the $\varepsilon$-active index sets $\{\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})\}_{s=1}^{N_{\nu}}$, respectively. The randomized selection combined with the sampling scheme enables the algorithm to solve subproblems of small sizes in the early iterations. By the convexity of the surrogate function, in step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ can be obtained via the state-of-the-art convex programming solvers. However, since the indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}$ are randomly selected from the $\varepsilon$-active index sets, the touching condition (P1) may not be satisfied when $\varepsilon>0$; namely, the surrogate function value $\widehat{F}_{\nu}(\theta^{\nu};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ might be strictly larger than $F_{\nu}(\theta^{\nu})$. Thus the candidate solution $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update with the desired descent property. Hence, in step 5 the candidate solution $\theta^{\nu+\frac{1}{2}}$ is accepted only when sufficient descent is satisfied. In the special case of $\varepsilon=0$ for which the touching condition (P1) is satisfied, the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ satisfies the sufficient descent inequality in step 5, and thus we could directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
@@
Next, we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (10). Due to the combination of sampling and enhanced technique, we introduce the averaged residual for convergence analysis, which is technically different from the analysis in Pang et al. (2016), Cui et al. (2018), Lu et al. (2019). We present the main theoretical results with essential discussions while leaving technical details in Online Appendix 8. Throughout the analysis, we make several blanket assumptions: the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$ and any $\bar{\theta}\in\Theta$.
@@
Theorem 3.6 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity. Moreover, under the aforementioned settings where $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ and a regularity assumption on $F$, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (10) with probability 1. This indicates that the PADR-based ERM model together with the ESMM algorithm is asymptotically optimal to the original problem (2). However, evidenced by the numerical implementations, when $\varepsilon$ is large, the candidate solutions produced in step 4 of Algorithm 1 may be rejected by the update rule in step 5 very often, which results in inefficient computation and slow convergence. Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy which will be specified in Section 4.
@@
The enhanced $\varepsilon$-active surrogation with positive $\varepsilon$ is essential for general cases. As noted in Liu et al. (2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity. This is also supported with Example 3.1 in Section 3.1.

### v5

@@ 2.3 Consistency comparison with existing DR-based methods
We compare the consistency result of the PADR-based ERM model with two prominent types of DR-based ERM methods in Table 2.3. NV-ERM1 and NV-ERM2 represent the two LDR methods from Ban and Rudin (2018), the latter of which additionally applies regularization in the objective. $\text{RKHS-DR}_{\text{NP}}$ and $\text{RKHS-DR}_{\text{BK}}$ represent RKHS-DR methods from Notz and Pibernik (2021) and Bertsimas and Koduri (2021) respectively, with the former imposing regularization via constraints and the latter applying regularization in the objective.
@@
Comparison of error bounds for multiple DR-based ERM models. Problem type Method Approximation error Estimation error Excess risk Convex NV-ERM1 0 (Linear asm.) $\widetilde{\mathcal{O}}(n^{-\frac{1}{2+p/2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2+p/2}})$ NV-ERM2 0 (Linear asm.) $\geq\mathcal{O}(n^{-\frac{1}{2+p/4}})$ / $\text{RKHS-DR}_{\text{NP}}$ / $\mathcal{O}(n^{-\frac{1}{2}})$ / $\text{RKHS-DR}_{\text{BK}}$ 0 $\geq\mathcal{O}(n^{-\frac{1}{2}})$ / Nonconvex PADR $\mathcal{O}(K^{-\frac{1}{p}})$ $\widetilde{\mathcal{O}}(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{4+p}})$ The complexity with $\geq$” means that the error is only bounded partly. The notation “/” means that the error bound is not provided.
Out of the four benchmarks considered in the literature, only Ban and Rudin (2018) provide an excess risk bound for NV-ERM1, leveraging the special structure of the Newsvendor problem with a linear true demand model. The estimation error bounds for NV-ERM2 and $\text{RKHS-DR}_{\text{BK}}$ are not provided due to the unknown regularization bias. Since an RKHS with a universal kernel is dense in the space of continuous functions, there is no approximation error for $\text{RKHS-DR}_{\text{BK}}$ , while such an error of $\text{RKHS-DR}_{\text{NP}}$ with a norm constraint is not analyzed in Notz and Pibernik (2021). It should be noted that the theoretical consistency results for all benchmarks in the aforementioned references, except $\text{RKHS-DR}_{\text{NP}}$ , are restricted to convex SP problems.
In comparison, our PADR-based ERM model is nonasymptotically consistent in more general cases, including convex SP problems and nonconvex ones. Moreover, by Corollary 2.4, if there exists a convex or DC optimal DR, our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large. This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $M,M]^{d}$ by employing a truncated PADR class ${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{p,d}(\mu)$ in Theorem 2.3 could be independent of $\mu$.
@@ 3 Efficient Sampling-based Algorithm for Learning PADR
where the objective is the finite sum of composite-PA functions. Instead of attempting to achieve global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions. Among various kinds of stationary concepts, the d-stationary solution is the sharpest type of first-order stationary solutions for general nonconvex optimization. For the PADR learning problem (9) with a general nonconvex and nondifferentiable composite structure, our algorithmic development is mainly motivated by the following question:
@@
Building on this motivation, we develop an MM-type algorithm specializing to the structured PADR-based ERM problem with the theoretical convergence analysis. It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which is up to 4 times faster than the basic MM method.
To make our discussion self-contained, we first present the standard MM algorithms in Section 3.1, which serves as a reference for the motivation behind our proposed sampling-based algorithm in Section 3.2. We make some remarks on notations to facilitate algorithmic development and convergence analysis. To emphasize the role of PADR as a function of $\theta$, with $\xi=(x,y)\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,x)$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,x)$ where $\{g_{i}(\theta,x)\}$ and $\{h_{i}(\theta,x)\}$ are linear functions with respect to $\theta$, and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
@@ 3.1 Preliminary analysis: standard MM algorithms
A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021a, Definition 7.1.1). Under (P1) and (P2), it yields the sequence of decreasing objective values as follows,
@@
Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by state-of-the-art convex-programming solvers. By the basic analysis (see Chapter 7 in Cui and Pang (2021a)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$. Such a fixed point property could be related to various types of stationarity of the original problem depending on the surrogate functions as illustrated in Cui et al. (2018a) and Liu et al. (2022a).
Regarding the PADR-based ERM problem (10), due to the nonconvex and nondifferentiable piecewise affine functions, a major issue of the standard MM algorithms including Lipp and Boyd (2016) and Le Thi and Pham Dinh (2018) is that the fixed point corresponding to the surrogation map is a (compound) critical solution (see definitions in Pang et al. (2016), Liu et al. (2022a)) which is much weaker than directional stationarity. Specifically, based on the composite structure in (10), approaches to construct surrogate functions of $F(\theta,\xi)$ in Cui et al. (2018a) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
@@
Utilizing the above upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$, the following example illustrates the construction of the surrogate function when outer function $\varphi(\cdot,\xi)$ is linear.
Example 3.1
For $\varphi(z,\xi)=c_{\xi}^{\top}z$, where $c_{\xi}\in\mathbb{R}^{d}$ is a parameter dependent on $\xi$, the composite objective function is $F(\theta,\xi)=c_{\xi}^{\top}f(\theta,\xi)$, and its surrogate function is constructed as
$\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{c_{\xi},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{c_{\xi},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1}),$
where $I\coloneqq\big(I_{1},I_{2}\big)\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$. For $d>1$, the maximum is taken componentwisely.
Surrogation in Example 3.1 obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm. Clearly $f^{\prime}(\cdot,\xi)(\theta;d)=g^{\prime}(\cdot,\xi)(\theta;d)-\max\{\,\nabla h_{i}(\bar{\theta},\xi)^{\top}d:\,i\in\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi)\}$, whereas the surrogate function $\widehat{F}(\theta,\xi;\bar{\theta},I)$ employs an arbitrary active index $I\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$. Hence, the fixed point of the surrogation map is a critical solution. Solutions of criticality are not desirable because in practice solutions may be at the “cliff” of the cost landscape, which is illustrated in Example 3.2 and Figure 1b.
@@
Consider problem (10) with $d=2$, $q=1$, $\Xi=\{\xi_{0}\}$, $\varphi(z,\xi_{0})=z_{1}^{2}+z_{2},f_{1}(\theta,\xi_{0})=\theta$, and $f_{2}(\theta,\xi_{0})=-\max\{-\theta,0\}$. In this case, the composite objective function is $F(\theta)=\theta^{2}+\max\{-\theta,0\}$. Figure 1a illustrates the two candidates of surrogate functions depending on which piece of the inner PA function $f_{2}(\theta,\xi_{0})$ is active at the reference point of interest. Figure 1b and Figure 1c illustrate the iterative procedures of the standard MM algorithm and the enhanced MM algorithm respectively, with the initial point $\theta_{0}=1.1$. The standard MM algorithm converges to a critical point $\theta_{\infty}=0$. For the enhanced MM algorithm with $\varepsilon=0.3$, at the second iterate point $\theta_{2}$, we have $\left\lvert\mathcal{I}^{\varepsilon}(\theta_{2},\xi_{0})\right\rvert=2$ at $\theta_{2}$, which thus leads to the next iterate point passing over 0 and continues to move towards a d-stationary point $\theta^{d}$ that is also globally optimal to the problem.
@@
Despite possessing stronger convergence properties, the enhanced MM algorithm applied for large-scale ERM problems could be exceptionally inefficient as it requires iteratively solving a potentially very large number of finite-sum convex subproblems over all the data. Sampling-based MM-type algorithms have been recently studied for (partially) smooth stochastic DC programs in Nitanda and Suzuki (2017), Xu et al. (2019), and Le Thi et al. (2022) and nondifferentiable stochastic composite DC problems in Liu et al. (2022a) to compute compound critical solutions. Based on a natural idea of combining a sequential sampling scheme with the enhanced technique, we propose an enhanced sampling-based MM algorithm with rigorous convergence analysis in the next subsection to solve the stochastic composite DC programs (10) with both computational efficiency and theoretical convergence guarantee.
@@ 3.2 Enhanced sampling-based MM algorithm for learning PADR
The directional derivative consistency condition is introduced in Cui and Pang (2021a) to relate the directional derivatives of surrogation and original functions. We consider the following three cases of the outer function $\varphi$ covering a broad class of cost functions.
@@
For the above cases, we can construct the surrogate function of $F(\cdot,\xi)$ satisfying Assumptions (B1) to (B4), by utilizing surrogate functions (12) and (13) of the inner PA function $f(\cdot,\xi)$ with the selected index $I=(I_{1},I_{2})$. The specific construction and verification are provided in Online Appendix 7. In particular, Case 2 is a multivariate extension of the one considered in Cui et al. (2018a). Case 3 is suitable for nonconvex smooth cost functions that are applicable for many practical problems.
@@
For a general nonconvex problem on a convex set with the directional differentiable objective, d-stationarity is identified as the sharpest kind among the first-order stationarity concepts. Specialized to the composite DC program (10), according to Cui et al. (2018a, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2. Lu et al. (2019) proposes the strengthened d-stationarity for structured DC programs which takes into account the $\epsilon$-active index set. We extend to the concept of composite $(\varepsilon,\rho)$-strong d-stationarity specialized for the PADR-based ERM problem (9). The composite strongly d-stationary point can be interpreted as the fixed point corresponding to the proximal mapping of the expected minimum surrogate function $\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};\bar{\theta})]$ based on the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfying Assumptions (B1)-(B4).
@@
Due to the requirement of a unique proximal mapping point in the defining condition, our concept is slightly stronger than $(\alpha,\eta)$-d-stationarity defined in Lu et al. (2019) when the problem is reduced to composite DC program. Moreover, it is weaker than composite $\varepsilon$-strong d-stationarity introduced in Qi et al. (2021b) (see also Cui and Pang (2021a, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis. Furthermore, in Online Appendix 8.1, we show that $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ under a regularity assumption on $F$. It implies that the definition of $\mathcal{S}^{d}_{\varepsilon,\rho}$ serves as a natural connection between d-stationarity and global optimality.
@@
We propose the ESMM algorithm with a natural idea of combining the randomized $\varepsilon$-active surrogation with sequential sampling presented in Algorithm 1. At iteration $\nu$, a subset of samples $\Xi_{\nu}=\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ of size $N_{\nu}$ is generated in step 2. Then in step 3 a sampling-based surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ is constructed which can be regarded as the upper surrogation function to the sample average approximation function $F_{\nu}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot,\widetilde{\xi}_{\nu}^{s})$ with the combination of indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}=\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ randomly selected from the $\varepsilon$-active index sets $\{\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})\}_{s=1}^{N_{\nu}}$, respectively. The randomized selection combined with the sampling scheme enables the algorithm to solve subproblems of small sizes in the early iterations. By the convexity of the surrogate function, in step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ can be obtained via the state-of-the-art convex programming solvers. However, since the indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}$ are randomly selected from the $\varepsilon$-active index sets, the touching condition (P1) may not be satisfied when $\varepsilon>0$; namely, the surrogate function value $\widehat{F}_{\nu}(\theta^{\nu};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ might be strictly larger than $F_{\nu}(\theta^{\nu})$. Thus the candidate solution $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update with the desired descent property. Hence, in step 5 the candidate solution $\theta^{\nu+\frac{1}{2}}$ is accepted only when sufficient descent is satisfied. In the special case of $\varepsilon=0$ for which the touching condition (P1) is satisfied, the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ satisfies the sufficient descent inequality in step 5, and thus we could directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$. We remark that under such an update scheme, since the size of $\varepsilon$-active index set grows exponentially along with the number of pieces of PADR, the iterate point would be more easily rejected for PADR with larger pieces.
@@
Next, we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (10). Due to the combination of sampling and enhanced technique, we introduce the averaged residual for convergence analysis, which is technically different from the analysis in Pang et al. (2016), Cui et al. (2018a), Lu et al. (2019). We present the main theoretical results with essential discussions while leaving technical details in Online Appendix 8. Throughout the analysis, we make several blanket assumptions: the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$ and any $\bar{\theta}\in\Theta$.
@@
Theorem 3.7 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity that is stronger than d-stationarity by Proposition 3.5. Moreover, Proposition 8.2 in Online Appendix 8.1 shows that $\mathcal{S}^{d}_{\varepsilon,\rho}$ coincides with $\mathcal{S}^{*}$ under a sufficiently large $\varepsilon$ and a regularity assumption on $F$. Consequently, any limit point of ${\widetilde{\theta}^{,t}}_{t}$ generated by the algorithm is a global minimizer of (10) with probability one. This indicates that the PADR-based ERM model together with the ESMM algorithm is asymptotically optimal to the original problem (2). However, evidenced by the numerical implementations, when $\varepsilon$ is large, the candidate solutions produced in step 4 of Algorithm 1 may be rejected by the update rule in step 5 very often, which results in inefficient computation and slow convergence. Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy which will be specified in Section 4.
@@
The enhanced $\varepsilon$-active surrogation with positive $\varepsilon$ is essential for general cases. As noted in Liu et al. (2022a), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity. This is also supported by Example 3.2 in Section 3.1.